OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {positive integers divisible by 3} Answer 21 30 14 3

OpenStudy (anonymous):

14..as 14/3 will give you a rational

OpenStudy (anonymous):

3 does not go into .....?

OpenStudy (anonymous):

wait im lost which is it?

OpenStudy (anonymous):

21 / 3 = 7 30 / 3 = 10 3 / 3 = 1 14 / 3 ??????????

OpenStudy (anonymous):

first off...all of the above are positive integers...but not all of them are divisible by 3...14 is not divisible by 3. the rest are

OpenStudy (anonymous):

Okay sorry. Just checking. Sorry. Im not good at 7th grade math Im helping my sister.

OpenStudy (anonymous):

no problem:)